A satellite originally moves in a circular orbit of radius R around the Earth. Suppose it is moved into a circular orbit of radius 4R. What does the force exerted on the satellite then become?
(a) 8 times larger
(b) 4 times larger
(c) one-half as large
(d) one eighth as large
1
Expert's answer
2021-02-22T10:29:22-0500
"F=\frac{GmM}{R^2}\\F'=\frac{GmM}{(4R)^2}=\frac{GmM}{16R^2}=\frac{F}{16}"